(B) Internet service providers (ISP) need to resolve customer problems as quickly as possible. The probability for a customer call regarding Internet service interruption is resolved within one hour is 70%. During one day, the company had 10 customer calls.              (10 marks)

a)    What is the probability that exactly four of the customers had their problems solved? (Round your answer to the nearest thousandth).

 

 


1
Expert's answer
2021-06-07T10:31:27-0400

"P(x=k)=C^k_np^kq^{n-k}"

We have:

"p=0.7,q=1-p=0.3,n=10"


"P(x=4)=C^4_{10}\cdot0.7^4\cdot0.3^6=\frac{10!}{6!4!}\cdot0.7^4\cdot0.3^6=210\cdot0.7^4\cdot0.3^6=0.037"